OpenStudy (anonymous):

EVALUATE THE EXPONENTIAL FUNCTION FOR THREE POSITIVE VALUES OF X,THREE NEGATIVE VALUES OF X,AND AT X=0.SHOW YOUR WORK.USE THE RESULTING ORDERED PAIRS TO PLOT THE GRAPH.STATE THE DOMAIN AND RANGE OF THE FUNCTIONATTACH GRAPH AS LISTED AND REQUIRED.F(X)=E^-X-1.CAN ANY ONE PLEASE HELP ME OUT IN THIS.PLEASE IT WOULD BE GREATFUL.

OpenStudy (dumbcow):

evaluate means substitute value for the variable so for x =0, e^-0-1 = 1/e

OpenStudy (anonymous):

use any three positive and negative values. For example use x = 1,2,3 and x = -1,-2,-3 and find f(x). your graph would be f(x) | | | | | |________________________________>x

OpenStudy (anonymous):

can you give me the whole solution please

OpenStudy (anonymous):

No, sorry, I can only show you how to do your problem.

OpenStudy (anonymous):

do we have to takef(x)=1/e

OpenStudy (anonymous):

i have posted